OpenStudy (anonymous):

Please Help! A parabola has vertex (-4,7) and passes through the point (-3,8). What is the equation in vertex form?

OpenStudy (anonymous):

Vertex formula: y=alx-hl+k 8=al-3+4l+7 8=a+7 a=1 y=lx+4l+7

OpenStudy (anonymous):

I could be wrong...

OpenStudy (anonymous):

I'm trying to figure it out. Vertex is the worst ever..